OpenStudy (anonymous):

Phosphorus-32 has a half-life of 14.3 days. How many days will it take for a sample of phosphorus-32 to decay to one-eighth its original amount?

OpenStudy (anonymous):

The rate of decay is a constant for the process of radioactive decay. This means that after one half life there is half as much, then after another there will be half of the half (one quarter), then after another half life there will be one eighth - you should be able to calculate the answer from this explanation, good luck,

OpenStudy (mayankdevnani):

Nt=1/8 No=1 t1/2=14,3 t=....? We can find 't' using basic formula, Nt/No=[1/2]^t/(t1/2) (1/8)/1=[1/2]^t/14,3 (1/8)=[1/2]^t/14,3 (1/2)^3=[1/2]^t/14,3 -----> with eliminating 1/2, 3=t/14.3 ---> t=14,3*3 t= 42.9 days.

OpenStudy (mayankdevnani):

@seabourn